OpenStudy (anonymous):

[3.04]Owen scored eighty-five points on a one hundred-point test and twenty-six on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

[3.04]Owen scored eighty-five points on a one hundred-point test and twenty-six on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

The fraction one-sixth, the decimal four tenths, the fraction one-half, the fraction three-fifths, the decimal seventy-five hundredths would be the list by values least to greatest 1/6 = .1667 .4 = .40 1/2 =.50 3/5 =.60 .75 = .75